### v2

@@ Appendix C An iBPGM for QROT
In this section, we briefly discuss how to employ an inexact Bregman proximal gradient method (iBPGM) with Sinkhorn’s algorithm as a subsolver for solving the QROT problem (4.1). We refer readers to [59, Section 5] for more details. Specifically, the iBPGM with the entropy kernel function for solving (4.1) can be given as follows: let $X^{0}>0$ and $\phi(X):=\sum_{ij}x_{ij}(\log x_{ij}-1)$, at the $k$-th iteration, compute
@@
where $M^{k}:=C+\lambda X^{k}-\mu_{k}\log X^{k}$. Note that problem (C.2) has the same form as the entropic regularized optimal transport problem and hence can be readily solved by the popular Sinkhorn’s algorithm; see [34, Section 4.2] for more details. Specifically, let $\Xi^{k}:=e^{-M^{k}/\mu_{k}}$. Then, given an initial positive vector $\bm{v}^{k,0}$, the iterative scheme is given by
@@ Appendix D Efficient implementation of ripALM for BPDN
When applying ripALM to the BPDN problem (4.9), a key component is the efficient implementation of the semismooth Newton (Ssn) method for solving the non-smooth equation (4.16). The main computational task in Ssn lies in computing the Newton direction, which requires solving the linear system (4.18). In the following, we describe how this system can be solved efficiently.
@@ Appendix E A dual ADMM for BPDN
$\min_{\bm{u},\,\bm{v},\,\bm{y}}$ $\quad\delta_{\mathcal{B}_{1}^{\infty}}(\bm{u})+\widehat{\kappa}\|\bm{v}\|-\bm{b}^{\top}\bm{y}$ (E.1)
$\mathrm{s.t.}$ $\quad D^{\top}\bm{y}-\bm{u}=0,~-\bm{y}-\bm{v}=0.$
@@ Appendix F Second-order cone programming reformulation for BPDN
$\min_{\bm{s},\,\bm{r}\in\mathbb{R}^{n}}$ $\quad\bm{1}^{\top}\bm{r}$
$\mathrm{s.t.}$ $\quad-\bm{r}\leq\bm{s}\leq\bm{r},\quad\bm{r}\geq\bm{0},$
$\quad\big(\widehat{\kappa},D\bm{s}-\bm{b}\big)\in\mathcal{Q}^{m+1}.$

### v3

@@ Appendix C An iBPGM for QROT
In this section, we briefly discuss how to employ an inexact Bregman proximal gradient method (iBPGM) with Sinkhorn’s algorithm as a subsolver for solving the QROT problem (4.1). We refer readers to [49, Section 5] for more details. Specifically, the iBPGM with the entropy kernel function for solving (4.1) can be given as follows: let $X^{0}>0$ and $\phi(X):=\sum_{ij}x_{ij}(\log x_{ij}-1)$, at the $k$-th iteration, compute
@@
where $M^{k}:=C+\lambda X^{k}-\mu_{k}\log X^{k}$. Note that problem (C.2) has the same form as the entropic regularized optimal transport problem and hence can be readily solved by the popular Sinkhorn’s algorithm; see [29, Section 4.2] for more details. Specifically, let $\Xi^{k}:=e^{-M^{k}/\mu_{k}}$. Then, given an initial positive vector $\bm{v}^{k,0}$, the iterative scheme is given by
@@ Appendix D Efficient implementation of ripALM for BPDN
When applying ripALM to the BPDN problem (4.9), a key component is the efficient implementation of the semismooth Newton (Ssn) method for solving the non-smooth equation (4.16). The main computational task in Ssn lies in computing the Newton direction, which requires solving the linear system (4.18). We next describe how to solve it efficiently.
@@ Appendix E A dual ADMM for BPDN
$\min_{\bm{u},\,\bm{v},\,\bm{y}}\Big\{\delta_{\mathcal{B}_{1}^{\infty}}(\bm{u})+\widehat{\kappa}\|\bm{v}\|-\bm{b}^{\top}\bm{y}\;\mid\;D^{\top}\bm{y}-\bm{u}=0,~-\bm{y}-\bm{v}=0\Big\}.$ (E.1)
@@ Appendix F Second-order cone programming reformulation for BPDN
$\min_{\bm{s},\,\bm{r}\in\mathbb{R}^{n}}\;\Big\{\bm{1}^{\top}\bm{r}\;\mid\;-\bm{r}\leq\bm{s}\leq\bm{r},\quad\bm{r}\geq\bm{0},\quad\big(\widehat{\kappa},D\bm{s}-\bm{b}\big)\in\mathcal{Q}^{m+1}\Big\}.$